find the roots using bisector method. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

3x^4-8x^3-37x^2+2x+40
1
Expert's answer
2020-10-27T20:06:28-0400

Solution. Find the roots using bisector method.

Find two points such that a < b and f(a)* f(b) < 0.

Find the midpoint of a and b, point m

If f(m) = 0 (m is root of the equation); else follow the next step

1) Divide the interval [a, b]

2) If f(m)*f(b) <0, let a =m

3) Else if f(m) *f(a), let b = m

Repeat steps until f(m) = 0 or relative error is below 0.05%.


"ɛ_r=|{\frac{x_{m,i+1}-x_{m,i}}{x_{m,i+1}}}|\times 100"

Let a=4 and b=6


As result first root is x1=5.

Let a=0 and b=2



As result second root is x2=1.

Let a=-2.5 and b=-1.5



As result third root is x3=-2.

Let a=-1.5 and b=-0.5



For the eleventh iteration approximate percent relative error is below 0.05%. Therefore the fourth root is


"x_4=-1.3335"

Answer. x1=5; x2=1; x3=-2; x4=-1.3335.